OpenStudy (anonymous):

using the slope formula, find the slope of the line through the given points.(2,-3)and(6,-3) hat is the slope of the line ?select the correct choice below and if necessary ,fill in the answer box to complete your choice .

OpenStudy (anonymous):

m = \[m = \frac{ rise }{ run } = \frac{ y_{2} - y_{1} }{ x_{2}-x_{1} }\] Just plug in the points into that and it will give you the slope.

OpenStudy (anonymous):

If you plug the values in you would get \[\frac{ -3 - (-3) }{ 6 - 2 }\] The slope in this case would be 0 because\[\frac{ 0 }{ 4 }\] when simplifying